# Validating a Machine Learning Approach to Identify Quenched Jets in Heavy-Ion Collisions

### v1

@@ 1 Introduction
In heavy ion collisions, interactions with the QGP medium modify the total energy and the internal structure of the jets as compared to proton-proton collisions. Jet suppression  CMSRAA; Aad:2012vca; Abelev:2013kqa; Adam:2015ewa depends on factors such as jet flavor CMS:2013qak; Aad_jetflavr2023; Aad_jetmass2023, path-length Atlasjetpathl; CMSjetpathl and fluctuations in medium density FluctQGP; FluctQGP2023. To understand how lost energy is redistributed, studies have focused on jet-hadron correlations, such as jet shapes Chatrchyan:2013kwa; ALICE:2019whv; ATLAS:2019pid, and fragmentation functions Chatrchyan:2014ava; Aaboud:2017bzv. Beyond these, jet substructure measurements, including groomed jet observables Sirunyan:2017bsd; STAR:2021kjt; ALargeIonColliderExperiment:2021mqf; CMSjetmg; aad2023measurement and energy-energy correlators CMS:2024ovv, offer insights into jet evolution from parton showering to hadronization. Jet substructures are sensitive to the QGP resolution scale and color coherence aad2023measurement, thereby providing new perspectives on the parton energy loss.
@@ 2 Simulated samples
Both medium and vacuum jet samples are embedded in a realistic heavy-ion background. Identical subtraction algorithm are used to prevent the neural network from learning uncorrelated background features that could mimic quenching effects Goncalves:2025fpf. The background simulation uses Angantyr Bierlich_2018 within Pythia 8.3 SciPostPhysCodeb.8 for 0–10% central heavy-ion events at $\sqrt{s_{\mathrm{NN}}}=$5.02\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$\$. The outgoing partons have a maximum transverse momentum of $\hat{p}_{\mathrm{T}}=$5\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$ [^1] For more information, see main42.cc and main42.cmnd in Pythia 8.3 examples.. In Figure 1, the average momentum density for each simulated underlying event is correlated with the multiplicity and is found to be consistent with CMS measurements Hayrapetyan:2897965. The average momentum density obtained here is consistent with that reported in our previous study  VandyMLJets, but the present method has the advantage of capturing event-by-event fluctuations. More details of the uncorrelated underlying events are listed in Table 1.
@@
We apply the machine learning approach to the photon-jet samples because the photon energy can be used to calibrate the jet energy loss independently. Similar to what we would do in experiments, we select Jewel events where the highest-$p_{T}$ photon and highest-$p_{T}$ jet in the event have an angular separation of $7\pi/8$. The jets selected for training (learning set) have back-to-back photons with $p_{T}\in[$200\text{\,}\mathrm{,}$$300\text{\,}\mathrm{]}$\text{~GeV}$, jet $p_{T}>$40\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$, and $|\eta|<2.0$. This photon energy cut during training constrains the neural network, ensuring that it learns primarily from energy loss rather than fragmentation bias. However, a well-trained model applied to jet quenching classification across different observables (classification set) can adapt to a photon energy range without an upper limit (e.g., $p_{T}^{\gamma}>$200\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$).
@@ 4 Machine learning setup
where $i,j$ denote the subjets at the declustering step $t$. In this way, sequential vectors $x_{0},...,x_{\mathrm{t}},$...] are extracted and used in the training of a neural network. Sequential vectors, which encode the showering history of the jet while preserving modifications from QGP medium, serve as input to train a neural network composed of LSTM and fully connected (FC) layers implemented in PyTorch NEURIPS2019_9015. As a type of recurrent neural network, an LSTM is particularly suited to learn from sequential data, using internal "gates" Sherstinsky_2020 to selectively retain or discard information over time steps. This allows it to effectively capture the patterns of jet-QGP interaction.
@@ 5.1 Training performance
The trained neural network can be used as a classifier to predict the probability that a jet is quenched. Figure 5 presents the distributions of the LSTM output and the receiver operating characteristic (ROC) curve. During training, unquenched jets are labeled as 0, while quenched jets are labeled as 1. The probability metric for jet quenching, learned during training, produces an output between 0 and 1 for each class.
@@ 5.2 Predictions for other observables
For the five quenching classes, we measure the distributions of the photon-jet energy fraction at the GEN level in Fig.7(a) and at the RECO level in Fig.7(b). With or without detector effects, the most quenched class (Q 0-20%) exhibits the greatest transverse momentum imbalance between the back-to-back jet and the photon. As the quenching level decreases toward the least quenched class (Q 80-100%), the transverse momentum between the jet and the photon becomes more balanced, with the ratio distribution shifting closer to zero (no energy loss). The detector effects broaden the distribution width in Fig. 7(b) compared to Fig. 7(a), but the ordering in the peak positions of the five quenching classes remains consistent.
@@
as shown in Fig. 8. At both the GEN level (Fig. 8(a)) and the RECO level (Fig. 8(b)), as the predicted quenching level increases, the jets exhibit greater enhancement in the large $\xi$ region, indicating that more soft particles enter the jet cone. Meanwhile, highly quenched jets (Q 0–20%) identified by ML show a greater depletion in the intermediate $\xi$ region compared to weakly quenched jets (Q 80-100%), indicating an increased energy loss. The jet fragmentation functions exhibit varying degrees of modification relative to the baseline, represented by the ratio of inclusive Jewel-Med jets to inclusive Jewel-Vac jets, categorized according to their ML-determined quenching levels. The RECO level results recover the GEN level ones with tracking efficiency corrected.
@@
The ML classification results for the jet momentum profile (integrated jet shape),
@@
are shown in Fig. 9. At both the GEN level (Fig. 9(a)) and the RECO level (Fig. 9(b)), the jets exhibit greater enhancement in the larger $\Delta R$ and greater depletion in the smaller $\Delta R$ as the predicted quenching level increases. This modification trend is consistent with the scenario in which strongly quenched jets exhibit a greater redistribution of jet energy to softer particles, extending to large angles away from the jet axis. Meanwhile, the least quenched jets (Q 80–100%) predicted by the neural network behave like vacuum jets, with their ratio to inclusive pp jets following a flatter trend compared to other quenched classes.
@@ 6 Conclusion
We demonstrated that a well-trained ML classifier based on sequential jet substructures can effectively distinguish true quenching features from other sources that may mimic quenching, such as thermal background and detector effects. To validate this, we cross-checked observables not included in the training process, e.g. photon-jet momentum imbalance, jet fragmentation function, and jet momentum profile. We found that the jets identified by the ML classifier with varying quenching levels exhibit corresponding modifications in these observables.
Using the Delphes simulation of the CMS detector response and particle flow candidate equivalents, we predicted the results of applying the ML method to data analysis. The results indicated that, even with detector effects, our ML classifier can still effectively learn from quenching features, consistent with jet energy loss estimated from the photon energy. We demonstrated significant potential of the machine learning approach for application to experimental data, where it can analyze a wide range of jet observables on a jet-by-jet basis. Such studies will help disentangle a variety of competing mechanisms that lead to jet modifications in QGP.

### v2

@@ 1 Introduction
In heavy ion collisions, interactions with the QGP medium modify the total energy and the internal structure of the jets as compared to proton-proton collisions. Jet suppression  CMSRAA; Aad:2012vca; Abelev:2013kqa; Adam:2015ewa depends on factors such as jet flavor CMS:2013qak; Aad_jetflavr2023; Aad_jetmass2023, path-length Atlasjetpathl; CMSjetpathl and fluctuations in medium density FluctQGP; FluctQGP2023. To understand how lost energy is redistributed, studies have focused on jet-hadron correlations, such as jet shapes Chatrchyan:2013kwa; ALICE:2019whv; ATLAS:2019pid, and fragmentation functions Chatrchyan:2014ava; Aaboud:2017bzv. Beyond these, jet substructure measurements, including groomed jet observables Sirunyan:2017bsd; STAR:2021kjt; ALargeIonColliderExperiment:2021mqf; CMSjetmg; aad2023measurement and energy-energy correlators CMS:2024ovv, offer insights into jet evolution from parton showering to hadronization. Jet substructures are sensitive to the QGP resolution scale and color coherence aad2023measurement, providing new perspectives on parton energy loss.
@@ 2 Simulated samples
Both medium and vacuum jet samples are embedded in a realistic heavy-ion background. The identical subtraction algorithm is used to prevent the neural network from learning uncorrelated background features that could mimic quenching effects Goncalves:2025fpf. The background simulation uses Angantyr Bierlich_2018 within Pythia 8.3 SciPostPhysCodeb.8 for 0–10% central heavy-ion events at $\sqrt{s_{\mathrm{NN}}}=$5.02\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$\$. The outgoing partons have a maximum transverse momentum of $\hat{p}_{\mathrm{T}}=$5\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$ [^1] For more information, see main42.cc and main42.cmnd in Pythia 8.3 examples.. In Figure 1, the average momentum density for each simulated underlying event is correlated with the multiplicity and is found to be consistent with the CMS measurements Hayrapetyan:2897965. The average momentum density obtained here is consistent with that reported in our previous study  VandyMLJets, but the present method has the advantage of capturing event-by-event fluctuations. More details of the uncorrelated underlying events are listed in Table 1.
@@
We apply the machine learning approach to the photon-jet samples because the photon energy can be used to calibrate the jet energy loss independently. Similarly to what we would do in experiments, we select Jewel events where the highest-$p_{T}$ photon and highest-$p_{T}$ jet in the event have an angular separation of $7\pi/8$. The jets selected for training (learning set) have back-to-back photons with $p_{T}\in[$200\text{\,}\mathrm{,}$$300\text{\,}\mathrm{]}$\text{~GeV}$, jet $p_{T}>$40\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$, and $|\eta|<2.0$. This photon energy cut during training constrains the neural network, ensuring that it learns primarily from energy loss rather than fragmentation bias. However, a well-trained model applied to jet quenching classification across different observables (classification set) can adapt to a photon energy range without an upper limit (e.g., $p_{T}^{\gamma}>$200\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$\$).
@@ 4 Machine learning setup
where $i,j$ denote the subjets in the declustering step $t$. In this way, sequential vectors $x_{0},...,x_{\mathrm{t}},$...] are extracted and used in the training of a neural network. Sequential vectors, which encode the showering history of the jet while preserving modifications from QGP medium, serve as input to train a neural network composed of LSTM and fully connected (FC) layers implemented in PyTorch NEURIPS2019_9015. As a type of recurrent neural network, an LSTM is particularly suited to learn from sequential data, using internal "gates" Sherstinsky_2020 to selectively retain or discard information over time steps. This allows it to effectively capture the patterns of jet-QGP interaction.
@@ 5.1 Training performance
The trained neural network can be used as a classifier to predict the probability that a jet is quenched. Figure 5 presents the distributions of the LSTM output and the receiver operating characteristic (ROC) curve. During training, the unquenched jets are labeled as 0, while the quenched jets are labeled as 1. The probability metric for jet quenching, learned during training, produces an output between 0 and 1 for each class.
@@ 5.2 Predictions for other observables
For the five quenching classes, we measure the distributions of the photon-jet energy fraction at the GEN level in Fig.7(a) and at the RECO level in Fig.7(b). With or without detector effects, the most quenched class (Q 0-20%) exhibits the greatest transverse momentum imbalance between the back-to-back jet and the photon. As the quenching level decreases toward the least quenched class (Q 80-100%), the transverse momentum between the jet and the photon becomes more balanced, with the ratio distribution shifting closer to zero (no energy loss). The ordering in the peak positions of the five quenching classes remains consistent when the detector effects are included.
@@
as shown in Fig. 8. At both the GEN level (Fig. 8(a)) and the RECO level (Fig. 8(b)), as the predicted quenching level increases, the jets exhibit greater enhancement in the large $\xi$ region, indicating that more soft particles enter the jet cone. Meanwhile, ML-identified high-quenched jets (Q 0–20%) show a greater depletion in the intermediate $\xi$ region compared to weakly quenched jets (Q 80-100%), indicating increased energy loss. The jet fragmentation functions exhibit varying degrees of modification relative to the baseline, represented by the ratio of inclusive Jewel-Med jets to inclusive Jewel-Vac jets, categorized according to their ML-determined quenching levels. The RECO level results recover the GEN level ones with tracking efficiency corrected.
@@
ML classification results for the jet momentum profile (integrated jet shape),
@@
are shown in Fig. 9. At both the GEN level (Fig. 9(a)) and the RECO level (Fig. 9(b)), the jets exhibit greater enhancement in the larger $\Delta R$ and greater depletion in the smaller $\Delta R$ as the predicted quenching level increases. This modification trend is consistent with the scenario in which strongly quenched jets exhibit a greater redistribution of jet energy to softer particles, extending to large angles away from the jet axis. Meanwhile, the least quenched jets (Q 80–100%) predicted by the neural network behave like vacuum jets, with their ratios to inclusive pp jets following a flatter trend compared to other quenched classes.
@@ 6 Conclusion
We demonstrated that a well-trained ML classifier based on sequential jet substructures can effectively distinguish true quenching features from other sources that may mimic quenching, such as thermal background and detector effects. To validate this, we cross-checked observables not included in the training process, e.g. photon-jet momentum imbalance, jet fragmentation function, and jet momentum profile. We found that the jets identified by the ML classifier with varying quenching levels exhibit the corresponding modifications in these observables.
Using the Delphes simulation of the CMS detector response and particle flow candidate equivalents, we predicted the results of applying the ML method to data analysis. The results indicated that, even with detector effects, our ML classifier can still effectively learn from quenching features, consistent with the jet energy loss estimated from the photon energy. We demonstrated significant potential of the machine learning approach for application to experimental data, where it can analyze a wide range of jet observables on a jet-by-jet basis. Such studies will help disentangle a variety of competing mechanisms that lead to jet modifications in QGP.